In ratios, how would you represent the following:
4 minutes to 1 hour
and
125 hits to 500 times at bat

Found 2 solutions by venugopalramana, stanbon:
Answer by venugopalramana(3286)

In ratios, how would you represent the following:
4 minutes to 1 hour
in general as a rule ratios are dimensionless.so if n.r is in minutes d.r should also be in minutes..hence=4 mts/1 hr=4mts/60 mts=1/15
and
125 hits to 500 times at bat
=125/500=1/4..ratios are usually simplified to get smallest possible integers in n.r.and d.r.

Answer by stanbon(48545)

In ratios, how would you represent the following:
4 minutes to 1 hour
4 min/ 60 min = 1/15 or 1:15
----------------------------------
125 hits to 500 times at bat
125 hits/500 times at bat = 1/4 or 1:4
Cheers,
Stan H.